# Monte Carlo simulations of relativistic shock breakout from a stellar wind

### v1

@@ 5.2 Comparison of Shock Widths
Another notable discrepancy between the simulation and the analytical model regarding shock width lies in its dependence on the shock Lorentz factor. As mentioned above, the analytical model predicts a scaling of $\Delta\tilde{\tau}\propto\Gamma_{u}$ for $f_{esc}\gg\Gamma_{u}^{-2}$. Qualitatively, the tendency for larger $\Gamma_{u}$ to result in a wider shock can be understood as a consequence of the Klein-Nishina suppression of the cross section, which becomes more significant for faster shocks. Indeed, for infinite shocks, a clear trend of increasing shock width with Lorentz factor, $\Delta\tilde{\tau}\propto\Gamma_{u}^{3}$, is found both in simulations (ILN20a) and in the analytical study (GNL18). While this scaling breaks down in the mildly relativistic regime, the simulations (ILN20a) still indicate a positive correlation between $\Delta\tilde{\tau}$ and $\Gamma_{u}$, at least down to $\Gamma_{u}=2$. In contrast, although the analytical model also predicts a positive dependence of shock width on Lorentz factor for finite shocks—specifically, $\Delta\tilde{\tau}\propto\Gamma_{u}$ when $f_{esc}\gg\Gamma_{u}^{-2}$—our current simulations of finite shocks ($f_{esc}>0$) exhibit the opposite trend: the shock width decreases with increasing $\Gamma_{u}$. [^11] Since the analytical model reproduces the numerical results quite well at $\Gamma_{u}=2$, a positive dependence of shock width on Lorentz factor ($\Delta\tilde{\tau}\propto\Gamma_{u}$) may still be valid in this mildly relativistic regime. However, the transition point at which the scaling changes from positive to inverse remains unclear, owing to the absence of simulations in the intermediate range between $\Gamma_{u}=2$ and $6$. For example, at $f_{esc}\simeq 0.4$, the shock width is found to be approximately $\Delta\tilde{\tau}\sim 7\times 10^{-3}$, $1\times 10^{-3}$, and $3\times 10^{-4}$ for $\Gamma_{u}=2$, $6$, and $10$, respectively, indicating an inverse correlation with the Lorentz factor. Thus, an important implication of our simulation results is that higher Lorentz factor shocks propagating in a wind-like environment tend to break out at larger radii (i.e., at smaller optical depths), contrary to previous expectation by GNL18 which finds opposite trend. We will revisit this point in more detail in Section 6.2. This discrepancy can again be understood as a consequence of the breakdown of the analytical model discussed above. Since higher Lorentz factors lead to the emergence of more prominent subshock-generated high-energy photons, deceleration via $\gamma-\gamma$ attenuation between counterstreaming photons becomes more efficient, resulting in faster deceleration.
@@ 6 Implications for the properties of relativistic shock breakout from a stellar wind
The discrepancy between our simulation results and the analytical model of GNL18 has important implications for shock breakout predictions based on that model. As noted earlier, the analytical model predicts the shock width to scale approximately as $\Delta\tilde{\tau}\sim(m_{e}/m_{p})\Gamma_{u}f_{esc}^{-1}$. On the other hand, our simulations show that for high-Lorentz-factor shocks ($\Gamma_{u}\gtrsim 6$), the optical depth $\Delta\tilde{\tau}$ significantly deviates from this scaling. By fitting the $\Delta\tilde{\tau}$$f_{esc}$ relation in our simulations with an analytical function of the form $\Delta\tilde{\tau}\propto\Gamma_{u}^{a}f_{esc}^{b}$, we obtain the best-fit scaling:
$\Delta\tilde{\tau}\approx 1.0\times 10^{-2}\Gamma_{u}^{-2}f_{esc}^{-1.6},\quad(6\lesssim\Gamma_{u}\lesssim 10).$ (2)
As described earlier, this scaling indicates a narrower shock width for a given $\Gamma_{u}$ and $f_{esc}$ than predicted by the analytical model, at least in the relativistic regime.
@@
$\tau_{bo}\approx 1.0\times 10^{-2}\Gamma_{bo}^{-2}\quad(6\lesssim\Gamma_{bo}\lesssim 10).$ (3)
where $\Gamma_{\rm bo}$ is the Lorentz factor at breakout. This is in stark contrast with the analytical prediction of $\tau_{bo,{\rm GNL18}}\approx(m_{e}/m_{p})\Gamma_{\rm bo}$, suggesting that our numerical results predict a much smaller breakout optical depth. The ratio between the two can be expressed as $\tau_{\rm bo}/\tau_{bo,{\rm GNL18}}=1.8\times 10^{-2}\Gamma_{bo,1}^{-3}$, where $\Gamma_{bo,1}=\Gamma_{bo}/10$, indicating that breakout occurs at a much larger radius in our model than in the analytical estimate for the same $\Gamma_{bo}$. In the following sections (§6.1 and §6.2), we revisit the breakout emission analysis originally presented in GNL18, adopting the revised shock width scaling up to $f_{esc}\approx 1$ for relativistic shocks ($\Gamma_{u}\gtrsim 6$), and highlight how the resulting observables differ from those predicted by the analytical model.
@@ 6.1 Closure relation
$E_{bo}=\frac{4\pi\Gamma_{bo}^{2}\tau_{bo}r_{bo}^{2}c^{2}}{\kappa}\approx 5.7\times 10^{46}r_{bo,13}^{2}\kappa_{0.2}^{-1}~{}{\rm erg},$ (6)
where $\kappa_{0.2}=\kappa/(0.2~{}{\rm cm}^{2}{g}^{-1})$. Notably, unlike the earlier prediction $E_{bo}\propto\Gamma_{bo}^{3}$, our result shows no explicit $\Gamma_{bo}$ dependence, reflecting the revised scaling $\tau_{bo}\propto\Gamma_{bo}^{-2}$.
@@
$E_{bo}\approx 5.7\times 10^{46}\kappa_{0.2}^{-1}\left(\frac{t_{bo}}{1.7\,\mathrm{s}}\right)^{2}\left(\frac{kT_{obs}}{1.7\,\mathrm{MeV}}\right)^{4}\quad\text{erg}.$ (7)
This differs from the previous prediction of $E_{bo}\propto T_{obs}^{7}$, reducing the temperature dependence to $E_{bo}\propto T_{obs}^{4}$. However, an important caveat of this rough closure relation given above is that it has only been validated within a narrow range of Lorentz factors, $6\lesssim\Gamma_{bo}\lesssim 10$, as our simulations are limited to $\Gamma_{u}=6$ and $10$. Extending the analysis to a broader range of $\Gamma_{u}$ is necessary to determine whether this scaling remains valid beyond this regime. On the other hand, for mildly relativistic shocks with Lorentz factors close to $\Gamma_{bo}\sim 2$, the agreement between the analytical model and numerical results at $\Gamma_{u}\sim 2$ suggests that the original closure relation recover its validity in a limited range, $62\lesssim\Gamma_{bo}<$ 6. Again, further analysis is required to precisely determine the range of validity, which is beyond the scope of this study.
@@ 6.2 Properties wind shock breakout based on a spherical explosion model
Following the prescription provided by GNL18, we consider a shock breaking out of a star characterized by a density profile of the form $\rho\propto(R_{*}-r)^{n}$, where $R_{*}$ is the stellar radius and $r$ is the distance from the stellar center. We further assume that the star is surrounded by an extended, wind-like medium with a density profile given by $\rho\propto r^{-2}$. For the case of $n=3$, the ejecta, after breaking out from the stellar surface and subsequently accelerating, exhibit a mass profile described by $m(\Gamma)\propto\Gamma^{-2.1}$, where $m(\Gamma)$ denotes the mass of the ejecta with Lorentz factors greater than $\Gamma$ (Johnson & McKee, 1971; Pan & Sari, 2006). Correspondingly, the energy of ejecta with Lorentz factors greater than $\Gamma$ can be expressed as $E(>\Gamma)\approx E_{0}(\Gamma/\Gamma_{0})^{-1.1}$, where $\Gamma_{0}$ and $E_{0}$ represent the Lorentz factor and the energy, respectively, of the breakout layer—defined as the leading edge of the shocked stellar matter that has a pair-unloaded optical depth of $\delta\tilde{\tau}\simeq 1$. Following Nakar & Sari (2012), the Lorentz factor of this breakout layer is estimated as: $\Gamma_{0}\approx 45E_{53}^{1.7}M_{ej,5}^{-1.2}R_{*,11}^{-0.95}$, where $E_{ex}=10^{53}E_{53}~{}\mathrm{erg}$ is the explosion energy, $M_{ej}=5M_{ej,5}M_{\odot}$ is the stellar ejecta mass, and $R_{*}=10^{11}R_{*,11}~{}\mathrm{cm}$ is the stellar radius. The corresponding energy is given by: $E_{0}=4\pi R_{*}^{2}\Gamma_{0}c^{2}\kappa^{-1}$. Here, we assume a constant opacity $\kappa$ for both the stellar and wind regions.
@@
$\Gamma_{sh}\approx 1.9E_{53}^{0.7}M_{ej,5}^{-0.49}R_{*,11}^{-0.39}\tau_{w*}^{-0.39}f_{esc}^{-0.31},$ (8)
$r_{sh}\approx 3.8\times 10^{13}E_{53}^{1.4}M_{ej,5}^{-0.99}R_{*,11}^{0.22}\tau_{w*}^{0.22}f_{esc}^{0.97}\quad\text{cm}.$ (9)
@@
$E_{bo}\approx 8.1\times 10^{47}\kappa_{0.2}^{-1}E_{53}^{2.8}M_{\rm ej,5}^{-2}R_{*,11}^{0.44}\tau_{w*}^{0.43}\quad\text{erg},$ (10)
$t_{bo}\approx 1.7\times 10^{2}R_{*,11}\tau_{w*}\quad\text{s}.$ (11)
@@
$kT_{obs}\approx 330E_{53}^{0.7}M_{ej,5}^{-0.49}R_{*,11}^{-0.39}\tau_{w*}^{-0.39}~{}{\rm keV}.$ (12)
A particularly notable result is that the breakout duration, $t_{bo}$, is independent of the explosion parameters and instead depends solely on the properties of the ambient medium—specifically, the stellar radius $R_{*}$ and the wind optical depth $\tau_{w*}$. This arises from the fact that, as mentioned above, a larger breakout Lorentz factor $\Gamma_{bo}$ also corresponds to a larger breakout radius $r_{bo}$, with the scaling relation $r_{bo}\propto\Gamma_{bo}^{2}$ holding for fixed values of $R_{*}$ and $\tau_{w*}$. This stands in contrast to the prediction by GNL18, in which the breakout duration exhibits strong sensitivity to explosion parameters: $t_{bo}\approx 1.6\times 10^{3}E_{53}^{-5.1}M_{\rm ej,5}^{3.6}R_{*,11}^{3.85}\tau_{w*}^{3.86}~{}\text{s}$.
To clearly demonstrate the difference between the previous analytical model and our current results at $\Gamma_{bo}\gtrsim 6$, let us consider a case where $E_{53}=5$, with all other parameters fixed to their fiducial values ($R_{*,11}=M_{ej,5}=\tau_{w*}=1$). In this case, our model yields $\Gamma_{bo}\approx 6$ ($kT_{obs}\approx 1~{}{\rm MeV}$), $E_{bo}\approx 7\times 10^{49}{\rm erg}$, and $t_{bo}\approx 170{\rm s}$. In contrast, the analytical model predicts $\Gamma_{bo}\approx 19$ ($kT_{obs}\approx 3.3~{}{\rm MeV}$), $E_{bo}\approx 2\times 10^{49}{\rm erg}$, and $t_{bo}\approx 0.42{\rm s}$. These results indicate that for an explosion energetic enough to drive a highly relativistic shock in an extended wind region, the breakout duration is significantly longer in our model compared to previous estimates, while the total emitted energy and observed temperature remain of the same order.
We note that realizing a relativistic breakout by reducing $M_{\rm ej}$, instead of increasing $E_{ex}$, yields the same outcome: while our model predicts a stable breakout duration, the analytical model suggests a much shorter one. It is also worth noting that the longer breakout duration predicted by the current study is favorable for observational prospects, enhancing detectability and enabling time-resolved spectral analysis. However, our simulations currently cover only a limited portion of the parameter space, specifically $\Gamma_{u}=6$ and $10$, and a finite range of $f_{esc}$. To rigorously assess the robustness of the breakout estimates and validate the broader applicability of the scaling relation presented in Equation (2), a more systematic exploration across a wider range of $\Gamma_{u}$ and $f_{esc}$ is necessary.
@@
$L_{bol}\approx\frac{f_{esc}E_{sh}}{t_{obs}}\propto t_{obs}^{-0.16},$ (13)
@@
Regarding breakouts occurring at lower velocities, specifically in the mildly relativistic regime ($62\lesssim\Gamma_{bo}<$ 6), the dependence of the shock width on $\Gamma_{u}$ remains unclear and likely exhibits non-monotonic behavior. This is primarily due to changes in the dominant deceleration mechanism within this range. In parallel, the scaling with $f_{esc}$ also varies significantly—shifting from $\Delta\tilde{\tau}\propto f_{esc}^{-1.6}$ at $\Gamma_{u}\approx 6$ to $\Delta\tilde{\tau}\propto f_{esc}^{-1}$ at $\Gamma_{u}\approx 2$, at least for $f_{esc}\gtrsim 0.1$. Consequently, it is challenging to derive consistent breakout estimates in this regime based on a single scaling relation, as was done above. Nevertheless, our results suggest that the shock width tends to increase with decreasing $\Gamma_{u}$, at least down to some Lorentz factor between $\Gamma_{u}=6$ and $2$ for finite shocks. This again implies a trend in which explosions producing lower $\Gamma_{bo}$ tend to break out at smaller radii. However, this trend may reverse at some point between $\Gamma_{bo}=6$ and $2$, given that the analytical model reproduces the shock structure at $\Gamma_{u}=2$ quite well. A detailed and systematic study covering this intermediate $\Gamma_{u}$ range is therefore needed to identify the transition point, which is beyond the scope of the current work.
@@
7 Summary and discussion

### v2

@@ 5.2 Comparison of Shock Widths
Another notable discrepancy between the simulation and the analytical model regarding shock width lies in its dependence on the shock Lorentz factor. As mentioned above, the analytical model predicts a scaling of $\Delta\tilde{\tau}\propto\Gamma_{u}$ for $f_{esc}\gg\Gamma_{u}^{-2}$. Qualitatively, the tendency for larger $\Gamma_{u}$ to result in a wider shock can be understood as a consequence of the Klein-Nishina suppression of the cross section, which becomes more significant for faster shocks. Indeed, for infinite shocks, a clear trend of increasing shock width with Lorentz factor, $\Delta\tilde{\tau}\propto\Gamma_{u}^{3}$, is found both in simulations (ILN20a) and in the analytical study (GNL18). While this scaling breaks down in the mildly relativistic regime, the simulations (ILN20a) still indicate a positive correlation between $\Delta\tilde{\tau}$ and $\Gamma_{u}$, at least down to $\Gamma_{u}=2$. In contrast, although the analytical model also predicts a positive dependence of shock width on Lorentz factor for finite shocks—specifically, $\Delta\tilde{\tau}\propto\Gamma_{u}$ when $f_{esc}\gg\Gamma_{u}^{-2}$—our current simulations of finite shocks ($f_{esc}>0$) exhibit the opposite trend: the shock width decreases with increasing $\Gamma_{u}$. [^11] Since the analytical model reproduces the numerical results well at $\Gamma_{u}=2$, a positive dependence of the shock width on Lorentz factor ($\Delta\tilde{\tau}\propto\Gamma_{u}$) may still hold within a quite narrow Lorentz-factor range around $\Gamma_{u}\simeq 2$ (for $\Gamma_{u}<3.5$). For example, for the simulations with highest $f_{esc}(=0.35-49)$, the shock width is found to be approximately $\Delta\tilde{\tau}\sim 6.7\times 10^{-3}$, $1.8\times 10^{-3}$, $7.6\times 10^{-4}$, $3.2\times 10^{-4}$, and $1.6\times 10^{-4}$ for $\Gamma_{u}=2$, $3.5$, 66, $10$, and $15$ respectively, indicating an inverse correlation with the Lorentz factor. Thus, an important implication of our simulation results is that higher Lorentz factor shocks propagating in a wind-like environment tend to break out at larger radii (i.e., at smaller optical depths), contrary to previous expectation by GNL18 which finds opposite trend. We will revisit this point in more detail in Section 6.2. This discrepancy can again be understood as a consequence of the breakdown of the analytical model discussed above. Since higher Lorentz factors lead to the emergence of more prominent subshock-generated high-energy photons, deceleration via $\gamma-\gamma$ attenuation between counterstreaming photons becomes more efficient, resulting in faster deceleration.
@@ 6 Implications for the properties of relativistic shock breakout from a stellar wind
The discrepancy between our simulation results and the analytical model of GNL18 has important implications for shock breakout predictions based on that model. As noted earlier, the analytical model predicts the shock width to scale approximately as $\Delta\tilde{\tau}\sim(m_{e}/m_{p})\Gamma_{u}f_{esc}^{-1}$. On the other hand, our simulations show that, for shocks with $\Gamma_{u}\gtrsim 3.5$, the optical depth $\Delta\tilde{\tau}$ departs from the analytical model. Focusing on the high-Lorentz-factor shocks ($\Gamma_{u}=6$, $10$, and $15$), where this departure is most evident, and restricting to the regime of appreciable escape ($f_{esc}\gtrsim 0.1$), we fit the simulation data with an analytical function of the form $\Delta\tilde{\tau}=C\,\Gamma_{u}^{a}f_{\rm esc}^{b}$ and obtain the following rough best-fit scaling:
$\Delta\tilde{\tau}\approx 4.0\times 10^{-3}\Gamma_{u}^{-1.8}f_{\rm esc}^{-1.9}\quad(6\lesssim\Gamma_{u}\lesssim 15).$ (2)
We do not include our $\Gamma_{u}=3.5$ simulations in this fit, because this Lorentz factor lies in a transitional regime where systematic deviations from the analytical model begin to emerge. Nevertheless, it is worth noting that the above scaling remains consistent with the $\Gamma_{u}=3.5$, $f_{esc}=0.49$ point. As described earlier, this scaling indicates a narrower shock width for a given $\Gamma_{u}$ and $f_{esc}$ than predicted by the analytical model, at least in the relativistic regime.
@@
$\tau_{bo}\approx 4.0\times 10^{-3}\Gamma_{bo}^{-1.8}\quad(6\lesssim\Gamma_{bo}\lesssim 15).$ (3)
where $\Gamma_{\rm bo}$ is the Lorentz factor at breakout. This is in stark contrast with the analytical prediction of $\tau_{bo,{\rm GNL18}}\approx(m_{e}/m_{p})\Gamma_{\rm bo}$, suggesting that our numerical results predict a much smaller breakout optical depth. The ratio between the two can be expressed as $\tau_{\rm bo}/\tau_{bo,{\rm GNL18}}=1.2\times 10^{-2}\Gamma_{bo,1}^{-2.8}$, where $\Gamma_{bo,1}=\Gamma_{bo}/10$, indicating that breakout occurs at a much larger radius in our model than in the analytical estimate for the same $\Gamma_{bo}$. In the following sections (§6.1 and §6.2), we revisit the breakout emission analysis originally presented in GNL18, adopting the revised shock width scaling up to $f_{esc}\approx 1$ for relativistic shocks ($\Gamma_{u}\gtrsim 6$), and highlight how the resulting observables differ from those predicted by the analytical model.
@@ 6.1 Closure relation
$E_{bo}=\frac{4\pi\Gamma_{bo}^{2}\tau_{bo}r_{bo}^{2}c^{2}}{\kappa}\approx 3.6\times 10^{46}r_{bo,13}^{2}\Gamma_{bo,1}^{0.2}\kappa_{0.2}^{-1}~{\rm erg},$ (6)
where $\kappa_{0.2}=\kappa/(0.2~{\rm cm}^{2}{g}^{-1})$. Notably, unlike the earlier prediction $E_{bo}\propto\Gamma_{bo}^{3}$, our result shows very weak dependence on $\Gamma_{bo}$, reflecting the revised scaling $\tau_{bo}\propto\Gamma_{bo}^{-1.8}$.
@@
$E_{bo}\approx 3.6\times 10^{46}\kappa_{0.2}^{-1}\left(\frac{t_{bo}}{1.7\,\mathrm{s}}\right)^{2}\left(\frac{kT_{obs}}{1.7\,\mathrm{MeV}}\right)^{4.2}\quad\text{erg}.$ (7)
This differs from the previous prediction of $E_{bo}\propto T_{obs}^{7}$, reducing the temperature dependence to $E_{bo}\propto T_{obs}^{4.2}$. However, an important caveat of this rough scaling relation is that it has only been examined over a limited range of Lorentz factor, $6\lesssim\Gamma_{\rm bo}\lesssim 15$. Extending the analysis to a broader range of $\Gamma_{u}$ is necessary to determine whether this scaling remains valid beyond this regime. On the other hand, for mildly relativistic shocks with Lorentz factors close to $\Gamma_{bo}\sim 2$, the agreement between the analytical model and numerical results at $\Gamma_{u}\sim 2$ suggests that the original closure relation recover its validity in a limited range, $2\lesssim\Gamma_{bo}<3.5$. Again, further analysis is required to precisely determine the range of validity, which is beyond the scope of this study.
@@ 6.2 Properties wind shock breakout based on a spherical explosion model
Following the prescription provided by GNL18, we consider a shock breaking out of a star characterized by a density profile of the form $\rho\propto(R_{*}-r)^{n}$, where $R_{*}$ is the stellar radius and $r$ is the distance from the stellar center. We further assume that the star is surrounded by an extended, wind-like medium with a density profile given by $\rho\propto r^{-2}$. For the case of $n=3$, the ejecta, after breaking out from the stellar surface and subsequently accelerating, exhibit a mass profile described by $m(\Gamma)\propto\Gamma^{-2.1}$, where $m(\Gamma)$ denotes the mass of the ejecta with Lorentz factors greater than $\Gamma$ (Johnson and McKee, 1971; Pan and Sari, 2006). Correspondingly, the energy of ejecta with Lorentz factors greater than $\Gamma$ can be expressed as $E(>\Gamma)\approx E_{0}(\Gamma/\Gamma_{0})^{-1.1}$, where $\Gamma_{0}$ and $E_{0}$ represent the Lorentz factor and the energy, respectively, of the breakout layer—defined as the leading edge of the shocked stellar matter that has a pair-unloaded optical depth of $\delta\tilde{\tau}\simeq 1$. Following Nakar and Sari (2012), the Lorentz factor of this breakout layer is estimated as: $\Gamma_{0}\approx 45E_{53}^{1.7}M_{ej,5}^{-1.2}R_{*,11}^{-0.95}$, where $E_{ex}=10^{53}E_{53}~\mathrm{erg}$ is the explosion energy, $M_{ej}=5M_{ej,5}M_{\odot}$ is the stellar ejecta mass, and $R_{*}=10^{11}R_{*,11}~\mathrm{cm}$ is the stellar radius. The corresponding energy is given by: $E_{0}=4\pi R_{*}^{2}\Gamma_{0}c^{2}\kappa^{-1}$. Here, we assume a constant opacity $\kappa$ for both the stellar and wind regions.
@@
$\Gamma_{sh}\approx 1.7E_{53}^{0.73}M_{ej,5}^{-0.51}R_{*,11}^{-0.41}\tau_{w*}^{-0.41}f_{esc}^{-0.39},$ (8)
$r_{sh}\approx 6.2\times 10^{13}E_{53}^{1.3}M_{ej,5}^{-0.93}R_{*,11}^{0.27}\tau_{w*}^{0.27}f_{esc}^{1.2}\quad\text{cm}.$ (9)
@@
$E_{bo}\approx 9.6\times 10^{47}\kappa_{0.2}^{-1}E_{53}^{2.8}M_{\rm ej,5}^{-2}R_{*,11}^{0.45}\tau_{w*}^{0.45}\quad\text{erg},$ (10)
$t_{bo}\approx 3.8\times 10^{2}E_{53}^{-0.15}M_{\rm ej,5}^{0.1}R_{*,11}^{1.1}\tau_{w*}^{1.1}\quad\text{s}.$ (11)
@@
$kT_{obs}\approx 280E_{53}^{0.73}M_{ej,5}^{-0.51}R_{*,11}^{-0.41}\tau_{w*}^{-0.41}~{\rm keV}.$ (12)
A particularly notable result is that the breakout duration, $t_{bo}$, is quite insensitive to the explosion parameters and instead predominantly determined by the properties of the ambient medium—specifically, the stellar radius $R_{*}$ and the wind optical depth $\tau_{w*}$. This arises from the fact that, as mentioned above, a larger breakout Lorentz factor $\Gamma_{bo}$ also corresponds to a larger breakout radius $r_{bo}$, with the $r_{bo}\propto\Gamma_{bo}^{1.8}$ holding for fixed values of $R_{*}$ and $\tau_{w*}$. This stands in contrast to the prediction by GNL18, in which the breakout duration exhibits strong sensitivity to explosion parameters: $t_{bo}\approx 1.6\times 10^{3}E_{53}^{-5.1}M_{\rm ej,5}^{3.6}R_{*,11}^{3.85}\tau_{w*}^{3.86}~\text{s}$.
To clearly demonstrate the difference between the previous analytical model and our current results at $\Gamma_{bo}\gtrsim 6$, let us consider a case where $E_{53}=5$, with all other parameters fixed to their fiducial values ($R_{*,11}=M_{ej,5}=\tau_{w*}=1$). In this case, our model yields $\Gamma_{bo}\approx 5.5$ ($kT_{obs}\approx 0.9~{\rm MeV}$), $E_{bo}\approx 9\times 10^{49}{\rm erg}$, and $t_{bo}\approx 300{\rm s}$. In contrast, the analytical model predicts $\Gamma_{bo}\approx 19$ ($kT_{obs}\approx 3.3~{\rm MeV}$), $E_{bo}\approx 2\times 10^{49}{\rm erg}$, and $t_{bo}\approx 0.42{\rm s}$. These results indicate that for an explosion energetic enough to drive a highly relativistic shock in an extended wind region, the breakout duration is significantly longer in our model compared to previous estimates, while the total emitted energy and observed temperature remain of the same order.
We note that realizing a relativistic breakout by reducing $M_{\rm ej}$, instead of increasing $E_{ex}$, yields the same outcome: while our model predicts a stable breakout duration, the analytical model suggests a much shorter one. It is also worth noting that the longer breakout duration predicted by the current study is favorable for observational prospects, enhancing detectability and enabling time-resolved spectral analysis. However, our simulations currently cover only a limited portion of the parameter space, specifically $\Gamma_{u}=6$, $10$, and $15$ and a finite range of $f_{esc}$. To rigorously assess the robustness of the breakout estimates and validate the broader applicability of the scaling relation presented in Equation (2), a more systematic exploration across a wider range of $\Gamma_{u}$ and $f_{esc}$ is necessary.
@@
$L_{bol}\approx\frac{f_{esc}E_{sh}}{t_{obs}}\propto t_{obs}^{-0.28},$ (13)
@@
Regarding breakouts occurring at lower velocities, specifically in the mildly relativistic regime ($2\lesssim\Gamma_{bo}<6$), the dependence of the shock width on $\Gamma_{u}$ remains unclear and likely exhibits non-monotonic behavior. This is primarily because the dominant deceleration mechanism changes within this range, as indicated by our intermediate $\Gamma_{u}=3.5$ simulations, which exhibit transitional properties. In parallel, the scaling with $f_{esc}$ also varies significantly—shifting from $\Delta\tilde{\tau}\propto f_{esc}^{-1.9}$ at $\Gamma_{u}\gtrsim 6$ to $\Delta\tilde{\tau}\propto f_{esc}^{-1}$ at $\Gamma_{u}\approx 2-3.5$, at least for $f_{esc}\gtrsim 0.1$. Consequently, it is challenging to derive consistent breakout estimates in this regime based on a single scaling relation, as was done above. Nevertheless, our results suggest that the shock width tends to increase with decreasing $\Gamma_{u}$, at least down to some Lorentz factor between $\Gamma_{u}=3.5$ and $2$ for finite shocks. This again implies a trend in which explosions producing lower $\Gamma_{bo}$ tend to break out at smaller radii. However, this trend may reverse at some point between $\Gamma_{bo}=3.5$ and $2$, given that the analytical model reproduces the shock structure at $\Gamma_{u}=2$ quite well. A detailed and systematic study covering this intermediate $\Gamma_{u}$ range is therefore needed to identify the transition point, which is beyond the scope of the current work.
@@
7 Limitations of the present analysis
Despite the valuable insights provided by our simulations, several limitations should be acknowledged. Here we briefly summarize these limitations.
7.1 On the assumption of steady-state and planar geometry
The key assumptions in the present study are that the shock structure can be described by a steady-state and planar geometry. For the two approximations to hold, an essentially identical requirement must be satisfied: the shock-crossing time, $t_{\rm cross}\sim\Gamma_{u}\Delta z/c$—the timescale for the shock to sweep up the mass contained in the shock transition layer of width $\Delta z$ measured in the shock rest frame—–must be much shorter than the dynamical timescale of the shock, which, in a wind-like environment, is given by $t_{\rm dyn}\sim r/c$; hence $t_{\rm cross}\ll t_{\rm dyn}$. While this condition is met well before the breakout, its validity becomes marginal during the breakout phase, as breakout in a wind-like environment leads to the two timescale to be comparable $t_{\rm cross}\sim t_{\rm dyn}$.
In this regard, recent work by Wasserman et al. (2025), based on dynamical simulations of sub-relativistic breakouts in a wind, argues that, contrary to the assumptions adopted in earlier studies (Ioka et al., 2019, ILN20b), the steady-state and planar approximations are likely to break down during breakout. We note, however, that relativistic and sub-relativistic breakouts have markedly different natures. In sub-relativistic RMS breakouts, once breakout begins it is essentially completed within a single dynamical time, i.e., $f_{\rm esc}$ evolves from 0 to 1 over $\sim t_{\rm dyn}$. By contrast, as discussed in GNL18, relativistic breakout proceeds much more gradually, spanning many dynamical times during the evolution from $f_{\rm esc}=0$ to 1. This difference arises because, in sub-relativistic RMSs, the shock optical depth cannot fall significantly below the diffusion length, i.e., $\Delta\tilde{\tau}\sim\beta_{u}^{-1}$, thereby limiting the breakout duration to $\sim t_{\rm dyn}$. In relativistic RMSs, however, self-generated opacity via pair production, allows the shock to propagate over several orders of magnitude in radius during breakout, as demontsrated in the present study.
More concretely, the scaling in Equation (2) indicates that, $f_{esc}\gtrsim 0.1$, the pair-unloaded optical depth of the shock scales as $\Delta\tilde{\tau}\propto f_{esc}^{-1.9}$. Combined with the optical depth scaling in a wind, $\tilde{\tau}\propto r^{-1}$, this implies $f_{esc}\propto r^{1/1.9}$. As a result, $f_{esc}$ increases by only a factor of $\approx 2^{1/1.9}\sim 1.4$ within one dynamical time, leading to a gradual breakout in which $f_{esc}$ can be treated as an adiabatic parameter. Thus, although the steady-state and planar assumptions are marginal in both sub-relativistic and relativistic breakouts, the two cases should be clearly distinguished. Accordingly, the conclusion of Wasserman et al. (2025) does not directly carry over to the relativistic case, and the much more gradual evolution makes the steady-state approximation at least more justifiable. Nevertheless, validating these assumptions ultimately requires dynamical simulations of relativistic shock breakouts—an important but challenging direction for future research.
7.2 On the effect of particle acceleration at the subshock
As demonstrated by our simulations, breakout from a wind is likely to be accompanied by a collisionless subshock whose strength increases as the escape fraction $f_{esc}$ rises. Because particle acceleration is expected to occur at the subshock, it may influence the high-energy spectrum; in particular, accelerated pairs could give rise to a non-thermal power-law extension in the emergent emission. However, if the energy carried by non-thermal pairs constitutes only a small fraction of that of the thermally heated pairs, their effect on the breakout spectra presented here is expected to be modest. This expectation arises because (i) the lower energy content of the non-thermal pairs implies a correspondingly smaller contribution of Comptonized photons than that produced by thermal pairs, and (ii) the higher-energy photons generated are subject to stronger $\gamma-\gamma$ opacity owing to the increased number density of target photons above the pair-creation threshold, thereby further suppressing the escaping flux. As a result, although non-thermal pairs may extend the spectrum toward higher energies, the associated flux is expected to remain relatively small. Moreover, it is worth noting that, owing to the small energy content in the counterstreaming non-thermal component, Comptonization by accelerated pairs does not affect the deceleration dynamics of the shock either. A quantitative prediction of the exact spectral shape would require a dedicated particle-acceleration model, which is beyond the scope of this work. At full breakout ($f_{\rm esc}=1$), when the entire shock has fully transitioned into a collisionless shock, non-thermal effects may become more relevant, but addressing this lies outside the scope of the present study. We defer a detailed investigation of this issue to future work.
7.3 On the decoupling among plasma particles
While we have adopted a single-fluid treatment of pairs and protons in the present study, the validity of this approximation remains uncertain. Specifically, as demonstrated by Levinson (2020), strong pair enrichment in RRMSs can lead to a decoupling between protons, electrons, and positrons, since the electric fields generated by small charge separations are no longer sufficient to maintain tight coupling. This decoupling may induce plasma instabilities and lead to microturbulence, as shown in Vanthieghem et al. (2022). Such plasma effects, if present, could modify the escaping spectra through inverse Compton scattering by high-energy pairs produced by the plasma effects and may also alter the deceleration process by changing the efficiency of $\gamma$-$\gamma$ pair production within the shock. Conversely, Mahlmann et al. (2023) showed that even a modest magnetic field, neglected here, may suffice to maintain tight coupling between species, potentially suppressing these instabilities. These issues warrant further investigation in future work.
8 Summary and conlusions